OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 5x – 9y = –16 2x + 6y = –16

OpenStudy (anonymous):

And it comes out to be 6x=32 right?

OpenStudy (lalaly):

sory typo:S u add both equations to eliminate y \[2 \times (5x-9y=-16)\]\[3 \times (2x+6y=-16)\] \[10x-18y=-32\]\[6x+18y=-48\]-------------- \[16x=-32\]

OpenStudy (anonymous):

Ohhh i missed the multiplying of the rest of the equation lol. Thank you :)

OpenStudy (anonymous):

16x = -80 though

OpenStudy (anonymous):

What do you mean?

OpenStudy (lalaly):

oh lols majre is right cos -48+-32 = -80

OpenStudy (anonymous):

Right!! My bad. Thank you, Majre lol

OpenStudy (anonymous):

after that you solve for x then input x into either equation to solve for y.